# Adaptive Reduced Tempering For Bayesian inverse problems and rare event simulation

### v1

@@ 4.2 Model #S: closed-form score
From an evaluation perspective, it is difficult to obtain closed-form reference posterior distributions in the context of the previous PDE model. In order to further evaluate the proposed algorithm with a closed-form characterizable reference probability distribution, we rely on the one-dimensional closed-form function proposed in [11]. In this toy model, the previous high-fidelity PDE solution $\Psi^{\ast}$ is substituted by a a one-dimensional function $\Psi^{\ast}:\mathbb{R}_{+}\to\mathbb{R}$, the RB approximation $\Psi$ is replaced by a cubic spline approximation $\Psi:\mathbb{R}_{+}\to\mathbb{R}$ and we assume that the error $|\Psi(x)-\Psi^{\ast}(x)|$ is directly available. For $x\in\mathcal{X}=\mathbb{R}_{+}$, the one-dimensional function is
@@
where $f(x)=15\left(\,{\rm l}{\rm l}_{x_{0}\leq x<x_{1}}\sin(x-x_{0})^{2}+\,{\rm l}{\rm l}_{x_{1}\leq x}(\sin(x_{1}-x_{0})^{2}-0.1(x-x_{1}))\right)$ with parameters $x_{0}=0.5$ and $x_{1}=5$.
@@ 4.3 Sampling problems
Figure 2: Problems #RE with model #RB. Illustration of the need for an adaptive approach to construct relevant reduced-score approximations of the PDE solution $\Psi^{\ast}$. The RB approximation $\Psi$ and its error $|\Psi-\Psi^{\ast}|$ are displayed for a sample $X$ in the rare event $\{x:\|\Psi^{\ast}(x)\|_{\infty}\geq 3\}$, at the start and end of the algorithm (notice the different error scales in the colorbar).
@@
Figure 3: Problem #RE with model #S. Illustration of the need for an adaptive approach to construct relevant approximations of $\Psi^{\ast}$ as defined in (4.1). Function $\Psi^{\ast}$ (blue curve), its spline approximation $\Psi$ (red curve) and the snapshots used to build it (green dots) are displayed at the start and end of the algorithm. The rare event is $\{x:\Psi^{\ast}(x)\geq 90\}$ (the “plateau” around $x=1e-2$).
@@ 4.3.1 Problem #RE: rare event simulation
In these numerical experiment, we consider a target distribution $\eta^{\ast}_{\beta_{\infty}}$ of the form (1.1) with $\mathrm{score}(\cdot)=1-\max(\ell-\|\cdot\|_{p},0)/\ell$ and $p>0$ yielding the score function
@@
Let $p^{*}=\pi(e^{\beta_{\infty}S^{\ast}})$ denote the target rare event probability. In the case of model #RB, this probability corresponds to the occurrence of reaching the critical temperature $\ell$, e.g. a melting temperature, on average in the domain ($p=1$) or in some point of the domain ($p=\infty$). For this model, we perform standard adaptive SMC simulation with a score based on the high-fidelity solution $\Psi^{\ast}(x)$ and using $1000$ particles. By averaging over 10 runs, we obtain after an intensive calculation empirical approximations of the target rare event probabilities: $p^{*}\approx 6.07e-04$ (for $p=1$ and $\ell=0.5$) and $p^{*}\approx 1.47e-04$ (for $p=\infty$ and $\ell=3$). In the case of model #S, the target probability $p^{*}\approx 2.18e-8$ (for $\ell=90$) is closed-form and analytically computable. We use the ART algorithm to estimate probability $p^{*}$, evaluating the unnormalized estimates $\hat{\gamma}^{N,H}_{\beta_{\infty},IS}(\,{\rm l}{\rm l})$ and $\hat{\gamma}^{N,H}_{\beta_{\infty},RSMC}(\,{\rm l}{\rm l})$ defined in (3.7) and (3.8).
@@
where we have used the definition of the score function, the triangular inequality, the Cauchy-Schwarz inequality, the Poincaré inequality with the constante $\pi^{-1}$ on the unit square domain and finally the a posteriori error inequality. As the finite element basis is an interpolating one, we will assume $h$ sufficiently large so that we can compute the $L^{1}$-norm as the average of the finite element coefficients. In the case of $p=\infty$, one can not in general rigorously derive an upper bound on the error relying on the a posteriori estimate $\Delta_{\Psi}(x)$. However, as the ART algorithm has the advantage of not requiring a rigorous error bound, we simply define the approximate error bound as
@@
For these low-temperature problems, ART performance is quantified by evaluating the relative expected square error as a function of the expected cost over $R$ independent runs. More specifically, let $\hat{p}_{r}$ be a short-hand notation for either the estimate $\hat{\gamma}^{N,H}_{\beta_{\infty},IS}(\,{\rm l}{\rm l})$ or $\hat{\gamma}^{N,H}_{\beta_{\infty},RSMC}(\,{\rm l}{\rm l})$ at the $r$-th run. After exhausting the budget $K$ of snapshots, the relative expected square error of the IS or RSMC estimator is approximated over runs by the empirical average
@@
and their expected cost (in units of the cost of evaluating $S^{\ast}$) is approximated by
$K+\frac{1}{R}\sum_{r=1,k=1}^{R,K}N^{(k)}_{r}\textrm{gain}^{(k)}$ (4.5)
with $\textrm{gain}^{(k)}$ representing the ratio of the cost of evaluating the reduced score $S^{(k)}$ to the cost of the true score $S^{\ast}$, and where $N^{(k)}_{r}$ represents the number of evaluations of the reduced score at the $k$-th iteration and $r$-th execution of the ART algorithm.
@@ 4.3.2 Problem #BI: Bayesian inverse problem
Figure 4: Problem #BI with model #RB. Illustration of the observations of the inverse problem. The high-fidelity solution $\Psi^{\ast}$ for a sample $X$ in the rare event set (obtained at the end of the algorithm for the sampling problem #RE) is observed on a random set of points up to Gaussian noise.
In this numerical experiment, we want to sample a Bayesian posterior distribution $\eta^{\ast}_{\beta_{\infty}}$ of the form (1.1) at the inverse temperature $\beta_{\infty}=1$, for $\pi$ being the log-normal prior distribution of model #RB, and given a set of observations following a Gaussian likelihood model. Let the supremum norm of the PDE solution and of its RB approximation evaluated at the diffusion parameter vector $x$ be denoted respectively by the functions $\varphi^{\ast}(\cdot)=\|\Psi^{\ast}(\cdot)\|_{\infty}$ and $\varphi(\cdot)=\|\Psi(\cdot)\|_{\infty}$. The objective of this Bayesian sampling problem is to estimate the posterior distribution of $X$, given the PDE solution on a finite set of points corrupted by noise. For evaluation purposes, we will focus in this experiment in the distribution of the scalar observables $\varphi^{\ast}(X)$ and $\varphi(X)$. In order to generate the observations, we consider the high-fidelity PDE solution $\Psi^{\ast}(x^{\star})$ evaluated at $x^{\star}$, where the latter parameter is sampled in the tail of the prior distribution $\pi$ (using a sample of the low-temperature simulation experiment, see Section 4.3.1). Let $\Psi^{\ast}_{z}\in\mathbb{R}$ be the PDE solution $\Psi^{\ast}$ taken at point $z\in\Omega$. Given a finite set of points $\mathcal{Z}\subset\Omega$, the set $\{\Psi^{\ast}_{z}(x^{\star})\}_{z\in\mathcal{Z}}$ is corrupted by an independent Gaussian noise of variance ${\sigma_{obs}^{2}}$ to generate the set of observations $\{\Psi^{obs}_{z}\}_{z\in\mathcal{Z}}$. The log likelihood of the Gaussian observation model is given by the score function
@@
In this experiment, we rely on the reduced score $S(x)=-\sum_{z\in\Omega_{obs}}(\Psi_{z}(x)-\Psi^{obs}_{z})^{2}/{\sigma_{obs}^{2}}$ where $\Psi_{z}(x)$ is the value of the RB approximation $\Psi(x)$ at the point $z\in\Omega$. Using (4.3), an approximate majoration is
@@
with the bound
$E(x)=\frac{1}{{\sigma_{obs}^{2}}}\sum_{z\in\Omega_{obs}}E_{\psi}(x)(E_{\psi}(x)+2|\Psi_{z}(x)-\Psi^{obs}_{z}|).$
@@
as a function of the expected cost (4.5).
@@ 4.4.1 Experimental setup for evaluation
ART is configured as follows for the different problems and models. The mutation proposal is based on an Ornstein-Uhlenbeck [19] adaptive state movement. The worst-case log cost threshold $c_{1}$, the adaptive tempering parameter $c_{2}$, the snapshot budget of $K$, the trigger parameter $j_{0}$, the number $M$ of Markov transitions at each mutation step, the number $R$ of runs and the number $N$ of particles are given, for each experiment, in the table below.
$c_{1}$ $c_{2}$ $K$ $j_{0}$ M $R$ $N\in$
problem #RE & model #RB 1e-2 1e-3 200 5 30 50 [1e2, 3e3]
problem #RE & model #S 1e-3 1e-3 200 5 100 100 [1e3, 1e4]
problem #BI & model #RB 1e-2 1e-3 200 5 30 100 [1e1, 5e2]
For problem #BI (which uses model #RB), we draw uniformly on the square $\Omega$ a set $\mathcal{Z}$ of $20$ random points $z$, where we observe the solution $\Psi_{z}^{\ast}$ up to a Gaussian noise of variance $21e-$ 2. The observations are displayed in Figure 4.
For problems#RB (resp. for problem #BI), we evaluate the empirical relative square error (4.4) (resp. the expected Kolmogorov-Smirnov statistic (4.6)) versus the average cost and compare to results obtained for several dozen of runs with the true score $S^{\ast}$ and for a varying number of particles, using either the standard AMS algorithm [8] (except for problem #BI, which can not be solved by AMS) or adaptive SMC simulation [3]. In the case of problem #RE, we also compare to performances obtained for several dozen of runs of the ARMS algorithm [11] with a varying number of particles.
@@ 4.4.2 Rare event simulation (Problem #RE)
Figure 5: Problem #RE (rare event simulation) & model #RB (PDE-based score). Relative expected square error as a function of the expected cost for ART’s IS and RSMC estimates, for a score function (4.2) with $p=1$ (left) and $p=\infty$ (right), and for a varying number of particles $N$. Comparison to the performance of adaptive tempering with $S^{\ast}$ and to results of AMS and ARMS computed in [11].
Figure 6: Problem #RE (rare event simulation) & model #S (closed-form score). Relative expected square error as a function of the expected cost for ART’s IS and RSMC estimates for a varying number of particles $N$. Comparison to the performance of adaptive tempering with $S^{\ast}$ and to results of AMS and ARMS computed in [11].
Figure 5 and  6 compare the performances of the ART algorithm with those of the state-of-the-art methods.
@@
Although ART’s RSMC and IS estimates are almost indistinguishable for model #RB, we note that ART’s RSMC estimate (which relies solely on the reduced score) deteriorates in the case of model #S, the square error tending to be constant with respect to the cost. For this case, using the spline approximation $S$ rather than the true score $S^{\ast}$ induces a significant bias. Indeed, the non-differentiability of $S^{\ast}$ at the rare event boundary leads to high-frequency oscillations (which tend to decrease with the number of interpolation points, but never disappear completely) in the spline model. The consequence is that a significant proportion of particles in the rare event set $\{x\in\mathcal{X}:S^{\ast}(x)\geq 1\}$ are to be excluded from the set $\{x\in\mathcal{X}:S(x)\geq 1\}$, leading to a negative bias. On the contrary, we observe that ART’s IS estimate is not biased and retains its good performance in this case, independently of the problem of convergence of $S$ to $S^{\ast}$.
For both standard and reduced estimates (ARMS and ART’s IS estimate), we observe approximately a unitary negative slope (on a logarithmic scale) of the variance as a function of cost, implying that the product of variance and cost is constant. Such a behavior suggests asymptotic normality of the proposed IS estimate, as it has been demonstrated for AMS [9].

### v2

@@ 4.2 Model #S: closed-form score
From an evaluation perspective, it is difficult to obtain closed-form reference posterior distributions in the context of the previous PDE model. In order to further evaluate the proposed algorithm with a closed-form characterizable reference probability distribution, we rely on the one-dimensional closed-form function proposed in [13]. In this toy model, the previous high-fidelity PDE solution $\Psi^{\ast}$ is substituted by a one-dimensional function $\Psi^{\ast}:\mathbb{R}_{+}\to\mathbb{R}$, the RB approximation $\Psi$ is replaced by a cubic spline approximation $\Psi:\mathbb{R}_{+}\to\mathbb{R}$ and we assume that the error $|\Psi(x)-\Psi^{\ast}(x)|$ is directly available. For $x\in\mathcal{X}=\mathbb{R}_{+}$, the one-dimensional function is
@@
where $f(x)=15\left(\,{\rm l}\hskip-1.99168pt{\rm l}_{x_{0}\leq x<x_{1}}\sin(x-x_{0})^{2}+\,{\rm l}\hskip-1.99168pt{\rm l}_{x_{1}\leq x}(\sin(x_{1}-x_{0})^{2}-0.1(x-x_{1}))\right)$ with parameters $x_{0}=0.5$ and $x_{1}=5$.
@@ 4.3 Sampling problems
Fig. 3: Problems #RE with model #RB. Illustration of the need for an adaptive approach to construct relevant reduced-score approximations of the PDE solution $\Psi^{\ast}$. The RB approximation $\Psi$ and its error $|\Psi-\Psi^{\ast}|$ are displayed for a sample $X$ in the rare event $\{x:\|\Psi^{\ast}(x)\|_{\infty}\geq 3\}$, at the start and the end of the algorithm (notice the different error scales in the colorbar).
@@
Fig. 4: Problem #RE with model #S. Illustration of the need for an adaptive approach to construct relevant approximations of $\Psi^{\ast}$ as defined in (4.1). Function $\Psi^{\ast}$ (blue curve), its spline approximation $\Psi$ (red curve) and the snapshots used to build it (green dots) are displayed at the start and end of the algorithm. The rare event is $\{x:\Psi^{\ast}(x)\geq 90\}$ (the “plateau” around x=1e2x=1e-2).
@@ 4.3.1 Problem #RE: rare event simulation
In these numerical experiments, we consider a target distribution $\eta^{\ast}_{\beta_{\infty}}$ of the form (1.1) with $\mathrm{score}(\cdot)=1-\max(\ell-\|\cdot\|_{p},0)/\ell$ and $p>0$ yielding the score function
@@
Let $p^{*}=\pi(e^{\beta_{\infty}S^{\ast}})$ denote the target rare event probability. In the case of model #RB, this probability corresponds to the occurrence of reaching the critical temperature $\ell$, e.g. a melting temperature, on average in the domain ($p=1$) or in some point of the domain ($p=\infty$). For this model, we perform standard adaptive SMC simulation with a score based on the high-fidelity solution $\Psi^{\ast}(x)$ and using $1000$ particles. By averaging over 10 runs, we obtain after an intensive calculation empirical approximations of the target rare event probabilities: $p^{*}\approx 6.07e-04$ (for $d=4$, $p=1$ and $\ell=0.5$), $p^{*}\approx 1.20e-04$ (for $d=25$, $p=1$ and $\ell=0.2$) and $p^{*}\approx 1.47e-04$ (for $d=4$, $p=\infty$ and $\ell=3$). In the case of model #S, the target probability $p^{*}\approx 2.18e-8$ (for $\ell=90$) is closed-form and analytically computable. We use the ART algorithm to estimate probability $p^{*}$, evaluating the unnormalized estimates $\hat{\gamma}^{N,H}_{\beta_{\infty},IS}(\,{\rm l}\hskip-1.99168pt{\rm l})$ and $\hat{\gamma}^{N,H}_{\beta_{\infty},RSMC}(\,{\rm l}\hskip-1.99168pt{\rm l})$ defined in (3.7) and (3.8).
@@
where we have used the definition of the score function, the triangle inequality, the Cauchy-Schwarz inequality, the Poincaré inequality with the constante $\pi^{-1}$ on the unit square domain and finally the a posteriori error inequality. As the finite element basis is an interpolating one, we will assume $h$ sufficiently large so that we can compute the $L^{1}$-norm as the average of the finite element coefficients. In the case of $p=\infty$, one cannot in general rigorously derive an upper bound on the error relying on the a posteriori estimate $\Delta_{\Psi}(x)$. However, as the ART algorithm has the advantage of not requiring a rigorous error bound, we simply define the approximate error bound as
@@
For these low-temperature problems, ART performance is quantified by evaluating the relative expected square error as a function of the expected cost over $R$ independent runs. More specifically, let $\hat{p}_{r}$ be a short-hand notation for either the estimate $\hat{\gamma}^{N,H}_{\beta_{\infty},IS}(\,{\rm l}\hskip-1.99168pt{\rm l})$ or $\hat{\gamma}^{N,H}_{\beta_{\infty},RSMC}(\,{\rm l}\hskip-1.99168pt{\rm l})$ at the $r$-th run. After exhausting the budget $K$ of snapshots, the relative expected square error of the IS or RSMC estimator is approximated over runs by the empirical average
@@
as a function of the expected computational cost (4.6) detailed below.
@@ 4.3.2 Problem #BI: Bayesian inverse problem
Fig. 5: Problem #BI with model #RB. Illustration of the observations of the inverse problem. The high-fidelity solution $\Psi^{\ast}$ for a sample $X$ in the rare event set (obtained at the end of the algorithm for the sampling problem #RE) is observed on a random set of points up to Gaussian noise.
In this numerical experiment, we want to sample a Bayesian posterior distribution $\eta^{\ast}_{\beta_{\infty}}$ of the form (1.1) at the inverse temperature $\beta_{\infty}=1$, for $\pi$ being the log-normal prior distribution of model #RB, and given a set of observations following some likelihood model. Let the supremum norm of the PDE solution and of its RB approximation evaluated at the diffusion parameter vector $x$ be denoted respectively by the functions $\varphi^{\ast}(\cdot)=\|\Psi^{\ast}(\cdot)\|_{\infty}$ and $\varphi(\cdot)=\|\Psi(\cdot)\|_{\infty}$. The objective of this Bayesian sampling problem is to estimate the posterior distribution of $X$, given the PDE solution on a finite set of points corrupted by noise. For evaluation purposes, we will focus in this experiment in the distribution of the scalar observables $\varphi^{\ast}(X)$ and $\varphi(X)$. In order to generate the observations, we consider the high-fidelity PDE solution $\Psi^{\ast}(x^{\star})$ evaluated at $x^{\star}$, where the latter parameter is sampled in the tail of the prior distribution $\pi$ (using a sample of the low-temperature simulation experiment, see Section 4.3.1). Let $\Psi^{\ast}_{z}\in\mathbb{R}$ be the PDE solution $\Psi^{\ast}$ taken at point $z\in\Omega$. Given a finite set of points $\mathcal{Z}\subset\Omega$, the set $\{\Psi^{\ast}_{z}(x^{\star})\}_{z\in\mathcal{Z}}$ is corrupted by an independent Gaussian noise of variance $\sigma_{obs}^{2}$} to generate the set of observations $\{\Psi^{obs}_{z}\}_{z\in\mathcal{Z}}$. The (non-Gaussian) log likelihood of the observation model is given by the score function
@@
In this experiment, we rely on the reduced score $S(x)=-\sum_{z\in\Omega_{obs}}(\Psi_{z}(x)-\Psi^{obs}_{z})^{2}/{\sigma_{obs}^{2}}$ where $\Psi_{z}(x)$ is the value of the RB approximation $\Psi(x)$ at the point $z\in\Omega$. Using (4.3), an approximate upper bound is
@@
with
$E(x)=\frac{1}{{\sigma_{obs}^{2}}}\sum_{z\in\Omega_{obs}}\Delta_{\Psi}(x)(\Delta_{\Psi}(x)+2|\Psi_{z}(x)-\Psi^{obs}_{z}|).$
@@
as a function of the expected computational cost (4.6) detailed below.
4.3.3 Expected Cost
In both problem #RE and #BI (Section 4.3.2 and 4.3.1) the expected cost (in units of the cost of evaluating $S^{\ast}$) is approximated by
$K+\frac{1}{R}\sum_{r=1,k=1}^{R,K}N^{(k)}_{r}\textrm{gain}^{(k)}$ (4.6)
with $\textrm{gain}^{(k)}$ representing the ratio of the cost of evaluating the reduced score $S^{(k)}$ and error $E^{(k)}$ to the cost of the true score $S^{\ast}$, and where $N^{(k)}_{r}$ represents the number of evaluations of the reduced score at the $k$-th iteration and $r$-th execution of the ART algorithm. Note that this expected cost neglects the extra-cost of updating the reduced model as compared to evaluating the true model. The gain is computed in practice empirically in an off-line phase. We provide in Appendix B more results for Model #RB using CPU time as a reference. For $d=4$, CPU time exhibits very mild discrepancy as compared to 4.6 and RB update is indeed computationally negligible. This is however no longer true as $d$ increases (around $d=25$).
@@ 4.4.1 Experimental setup for evaluation
ART is configured as follows for the different problems and models. The mutation proposal is based on an Ornstein-Uhlenbeck [23] adaptive state movement.
For each experiment, we give in the table below: the entropic log cost threshold $c_{1}$ triggering snapshot sampling, the adaptive tempering parameter $c_{2}$, the total number of snapshots $K$, the parameter $j_{0}$ triggering non-penalized IS of snapshots, the number $M$ of Markov transitions at each mutation step, the number $R$ of runs and the number $N$ of particles.
$h$ $d$ $K$ $j_{0}$ $M$ $R$ $N\in$ $c_{1}$ $c_{2}$
problem #RE & model #RB 5101 4 200 5 30 50 [1e2, 3e3] 1e-3 1e-2
problem #RE & model #RB 5101 25 200 50 30 20 {1e2} 1e-1 1e-2
problem #RE & model #S 1 1 200 5 100 100 [1e3, 1e4] 1e-3 1e-3
problem #BI & model #RB 5101 4 200 5 30 100 [1e1, 5e2] 1e-3 1e-2
For problem #BI (which uses model #RB), we draw uniformly on the square $\Omega$ a set $\mathcal{Z}$ of $20$ random points $z$, where we observe the solution $\Psi_{z}^{\ast}$ up to a Gaussian noise of variance 1e21e-2. The observations are displayed in Figure 5.
For problems #RE (resp. for problem #BI), we evaluate the empirical relative square error (4.4) (resp. the expected Kolmogorov-Smirnov statistic (4.5)) versus the average cost and compare to results obtained for several dozen of runs with the true score $S^{\ast}$ and for a varying number of particles, using either the standard adaptive multilevel splitting (AMS) algorithm [10] (except for problem #BI, which cannot be solved by AMS) or adaptive SMC simulation [5]. In the case of problem #RE, we also compare to the performance for several dozen of runs of the adaptive reduced multilevel splitting (ARMS) algorithm [13] with a varying number of particles. Under our experimental protocol, the number of evaluation of the true model in the reference multilevel splitting algorithms (AMS, ARMS) is typically of order $N\times 10^{3}$. For $N=100$, this yields around $500$ times more than in the ART case.
@@ 4.4.2 Rare event simulation (Problem #RE)
Fig. 6: Problem #RE (rare event simulation) & model #RB (PDE-based score, $d=4$). Relative expected square error as a function of the expected cost for ART’s IS and RSMC estimates, for a score function (4.2) with $p=1$ (left) and $p=\infty$ (right), and for a varying number of particles NN. Comparison to the performance of adaptive tempering with $S^{\ast}$ and to results of AMS and ARMS computed in [13].
Fig. 7: Problem #RE (rare event simulation) & model #S (closed-form score). Relative expected square error as a function of the expected cost for ART’s IS and RSMC estimates for a varying number of particles NN. Comparison to the performance of adaptive tempering with $S^{\ast}$ and to results of AMS and ARMS computed in [13].
Figure 6 and  7 compare the performances of the ART algorithm with those of the state-of-the-art methods in the case of small dimension of the parameter space: model #RB with $d=4$ or the one-dimensional model #S.
@@
Although ART’s RSMC and IS estimates are almost indistinguishable for model #RB, we note that ART’s RSMC estimate (which relies solely on the reduced score) deteriorates in the case of model #S, the square error tending to be constant with respect to the cost. For this case, using the spline approximation $S$ rather than the true score $S^{\ast}$ induces a significant bias. Indeed, the non-differentiability of $S^{\ast}$ at the rare event boundary leads to high-frequency oscillations (which tend to decrease with the number of interpolation points, but never disappear completely) in the spline model. The consequence is that a significant proportion of particles in the rare event set $\{x\in\mathcal{X}:S^{\ast}(x)\geq 1\}$ is excluded from the set $\{x\in\mathcal{X}:S(x)\geq 1\}$, leading to a negative bias. On the contrary, we observe that ART’s IS estimate is not biased and retains its good performance in this case, independently of the problem of convergence of $S$ to $S^{\ast}$.
On the other hand for both reference estimates (AMS and tempering) and IS estimates for ARMS and ART, we observe approximately a unitary negative slope (on a logarithmic scale) of the variance as a function of cost, implying that the product of variance and cost is constant. Such a behavior suggests asymptotic normality of the proposed IS estimate, as it has been demonstrated for AMS [11].